# Viscosity of polymer melts using non-affine theory based on vibrational modes

### v1

@@ IV SUMMARY AND CONCLUSIONS
We investigated the microscopic origin of viscosity in supercooled liquids from first principles using non-affine lattice dynamics (NALD). Using a model system of coarse-grained polymer melt (Kremer-Grest), we generated equilibrium configurations in the supercooled regime. These configurations were then used to calculate the vibrational density of states and affine force field correlators, which were subsequently used to compute the complex shear modulus. The viscosity of the system was directly obtained from the zero-frequency limit of the loss modulus. The temperature dependence of viscosity in the supercooled regime is accurately captured by this approach. For comparison, we also calculated the viscosity using the GK method and non-equilibrium shear MD simulations over the accessible temperature range. By comparing the viscosity results obtained from different methods in both the liquid and supercooled regimes, we demonstrated that the NALD approach reliably predicts the viscosity, particularly in the supercooled regime closer to glass transition where experimental measurements and simulations based on GK are challenging.
@@
The NALD framework requires a molecular friction parameter $\nu$ as an input. In this study, we have used a Langevin thermostat in the simulations with a fixed dissipation timescale $\tau=1$, which works as an effective inverse friction parameter for our system, and as such it was implemented in NALD, thus leaving no adjustable parameters. In a previous study by Palyulin et al. [59] it was shown that the viscoelastic response under the NALD framework could be studied as a Markovian system where the friction of the system can be controlled using the Langevin thermostat. In general, the friction kernel can be derived using a particle-bath Hamiltonian framework by solving the Euler-Lagrange equations for the coupled dynamics of a tagged particle and heat bath oscillators, following Zwanzig’s formalism [84, 85, 48]. In the NALD framework used thus far, the friction is a constant [45, 59, 60, 86]: in coarse-grained systems it matches the value of the Langevin thermostat, while in atomistic systems it is determined by matching the NALD calculation with the NEMD of the viscoelastic modulus at the high-frequency plateau. However, in experimental and real-world systems, the microscopic friction is typically time-dependent and non-Markovian [87, 75, 88, 89]. Although the memory kernel can be extracted from MD simulations [90, 91], a direct estimation based on the fluctuation-dissipation theorem [84, 85], associated with the generalized Langevin equation, can be obtained from the time autocorrelation function of the stochastic force [92]. Another approach to measure the memory kernel involves computing the momentum autocorrelation and force correlation functions [93]. Through the analysis of memory kernels, it should be possible to recover the expected long-time transport behavior. Accurate evaluation of the memory kernel under favorable conditions can provide insights into the viscoelastic response, facilitating the design of functional materials for practical applications [75, 60]. Hence, in future studies, the friction parameter should be handled as a parameter that is derived from the microscopic input of the system. Furthermore, this framework can be extended to investigate how different frequency modes in the vibrational spectrum (vDOS) influence the viscoelastic modulus and transport properties in atomistic systems [60], soft jammed amorphous [68], colloidal glass [94, 95], and granular fluids [96].
@@ Acknowledgments
AS, VV, and AZ gratefully acknowledge funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. VV acknowledges the computational resource provided via the project “TPLAMECH” on INDACO platform at the HPC facility of Università degli Studi di Milano. AZ gratefully acknowledges funding from the US Army Research Office through contract nr. W911NF-22-2-0256.

### v3

@@ IV SUMMARY AND CONCLUSIONS
We investigated the microscopic origin of viscosity in model supercooled polymeric liquids from first principles using non-affine lattice dynamics (NALD). Using a model system of coarse-grained polymer melt (Kremer-Grest), we generated equilibrium configurations in the supercooled regime. These configurations were then used to calculate the vibrational density of states and affine force-field correlators Zaccone (2023a), which were subsequently used to compute the complex shear modulus Zaccone (2023b). The viscosity of the system was directly obtained from the zero-frequency limit of the loss modulus. The temperature dependence of viscosity in the supercooled regime is accurately captured by this approach. For comparison, we also calculated the viscosity using the Green-Kubo (GK) method and non-equilibrium MD (NEMD) shear simulations over the accessible temperature range. By comparing the viscosity results obtained from different methods in both the liquid and supercooled regimes, we demonstrated that the NALD approach reliably predicts the viscosity, particularly in the supercooled regime closer to glass transition where experimental measurements and traditional simulation methods based on GK are challenging.
@@
The NALD framework requires a molecular friction parameter $\nu$ as an input. In this study, we have used a Langevin thermostat in the MD simulations with a fixed dissipation timescale $\tau=1$, which works as an effective inverse friction parameter for our system, and as such it was implemented in NALD, thus leaving no adjustable parameters. In a previous study by Palyulin et al. Palyulin et al. (2018) it was shown that the viscoelastic response in the NALD framework could be studied as a Markovian system where the friction of the system can be controlled using the Langevin thermostat. In general, the friction kernel can be derived using a particle-bath Hamiltonian framework by solving the Euler-Lagrange equations for the coupled dynamics of a tagged particle and heat bath oscillators, following Zwanzig’s formalism Zwanzig and Mountain (1965); Zwanzig (1973); Zaccone (2023b); Rognoni et al. (2021); Gottwald et al. (2016). In the NALD framework used thus far, the friction is a constant Lemaître and Maloney (2006); Palyulin et al. (2018); Vaibhav et al. (2024); Elder et al. (2019), because in coarse-grained systems it matches the value of the Langevin thermostat damping parameter. In the atomistic systems studied thus far, it is determined by matching the NALD calculation with the nonequilibrium molecular dynamics (NEMD) of the viscoelastic modulus at the high-frequency plateau Elder et al. (2019); Vaibhav et al. (2024). However, in experimental and real-world systems, the microscopic friction is typically time-dependent and non-Markovian Lickert and Stock (2020); Milster et al. (2024); Kowalik et al. (2019); Kiefer et al. (2025). Although the memory kernel can be extracted from MD simulations Schilling (2022); Scalfi et al. (2023), a direct estimation based on the fluctuation-dissipation theorem Zwanzig and Mountain (1965); Zwanzig (1973), associated with the generalized Langevin equation, can be obtained from the time autocorrelation function of the stochastic force Jung et al. (2017). Another approach to measure the memory kernel involves computing the momentum autocorrelation and force correlation functions Gottwald et al. (2016). Through the analysis of memory kernels, it should be possible to recover the expected long-time transport behavior. Accurate evaluation of the memory kernel under favorable conditions can provide insights into the viscoelastic response, facilitating the design of functional materials for practical applications Milster et al. (2024); Vaibhav et al. (2024). Hence, in future studies, the friction parameter should be handled as a parameter that is derived from the microscopic input of the system. Furthermore, this framework can be extended to investigate how different frequency modes in the vibrational spectrum (vDOS) influence the viscoelastic modulus and transport properties in atomistic systems Vaibhav et al. (2024), soft jammed amorphous Hara et al. (2025), colloidal glass Mason and Weitz (1995); Negi et al. (2014), polymer nanocomposites Mermet-Guyennet et al. (2015) and granular fluids Cheng et al. (2014). A crucial point for improvement in future work will be the scaling up of the eigenmodes computation towards lower frequencies, which is especially challenging for atomistic systems.
@@ Acknowledgments
AS, VV, and AZ gratefully acknowledge funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. VV acknowledges the computational resource provided via the project “TPLAMECH” on INDACO platform at the HPC facility of Università degli Studi di Milano. AZ gratefully acknowledges funding from the US Army DEVCOM Army Research Office through contract nr. W911NF-22-2-0256.